nightshade0773:

Belltown Library just added a new reading room with comfortable chairs so kids can relax with their books. In the center of the room, there is a rectangular rug that looks like a world map. The width of the rug is 12.5 feet, and the entire area is 200 square feet. Let ℓ represent the length of the rug. Which equation models the problem? Solve this equation to find the length of the rug. feet

crispyrat:

So height x width= area you have the width and the area so you have height*12.5=200 divide both sides by 12.5 height*(12.5/12.5)=200/12.5 height*1=200/12.5 so what is height

jhonyy9:

@crispyrat wrote:
So height x width= area you have the width and the area so you have height*12.5=200 divide both sides by 12.5 height*(12.5/12.5)=200/12.5 height*1=200/12.5 so what is height
you mean height wan be the length

jhonyy9:

bc. nowhere in the text of this problem not talk about height please more attention !!!

nightshade0773:

thats what got me

nightshade0773:

@crispyrat wrote:
So height x width= area you have the width and the area so you have height*12.5=200 divide both sides by 12.5 height*(12.5/12.5)=200/12.5 height*1=200/12.5 so what is height
it doesn't say anything about hieght so i guessed and got it coreect

jhonyy9:

@nightshade0773 wrote:
@crispyrat wrote:
So height x width= area you have the width and the area so you have height*12.5=200 divide both sides by 12.5 height*(12.5/12.5)=200/12.5 height*1=200/12.5 so what is height
it doesn't say anything about hieght so i guessed and got it coreect
yes correct result but for length not height